OpenStudy (cutiecomittee123):

if cscx=2 and x is in quadrant 2 then sin(2x)=

OpenStudy (nincompoop):

what is csc ?

OpenStudy (nincompoop):

I need to know what you know so I can help you properly

OpenStudy (cutiecomittee123):

csc = 2 csc= 1/sin uhh I dont really have any other info

OpenStudy (cutiecomittee123):

oh and it would be hypotenuse/opposite

OpenStudy (nincompoop):

okay, csc is the inverse of sin like you said \(csc = \large \frac{1}{sin} \)

OpenStudy (cutiecomittee123):

Yes

OpenStudy (nincompoop):

so it is possible to think it like \(\large \frac{1}{sin~x} = 2 \)

OpenStudy (nincompoop):

so if that is possible, it is also possible to think of other trigonometric identities. yes?

OpenStudy (cutiecomittee123):

Yes I think so

OpenStudy (nincompoop):

so tell me, what other trigonometric identities do you know or you're familiar with

OpenStudy (cutiecomittee123):

sin(2x)=2sinxcosx and I think that is important to know in this equation

OpenStudy (nincompoop):

I used the wrong word. I said inverse, when I meant to use RECIPROCAL.

OpenStudy (nincompoop):

YES! you can use something like that

OpenStudy (cutiecomittee123):

its okay same thing basically? Okay well I am not seeing how it can be used, I just though it would fit since we have the sin(2x)=

OpenStudy (cutiecomittee123):

so basically cscx=1/2sinxcox

OpenStudy (cutiecomittee123):

which leads me to thinking that sin(2x)=1/2 when cscx=2

OpenStudy (cutiecomittee123):

cosx*

OpenStudy (nincompoop):

does the word quadrant ring any bell?

OpenStudy (cutiecomittee123):

yes, and in the second quadrant- isnt everything positive?

OpenStudy (nincompoop):

hmm?

OpenStudy (nincompoop):



OpenStudy (cutiecomittee123):

Oh, wait sin is positive and cos is negative

OpenStudy (cutiecomittee123):

so then sin(2x)= -1/2 ??

OpenStudy (nincompoop):

don't be confused okay? isn't sin your y also cos is x this is why tan = sin/cos = y/x = slope

OpenStudy (cutiecomittee123):

Yes I know this.

OpenStudy (cutiecomittee123):

so csc=1/x

OpenStudy (nincompoop):

are you thinking like this? 1/x = 2 x = 1/2 so \(sin ~(2 \large \frac{1}{2}) = \)

OpenStudy (cutiecomittee123):

ohhh i see. I didn't think that far ahead. Thanks then. So now we just do (2)(1/2) which is 1

OpenStudy (nincompoop):

sorry my connection keeps lagging

OpenStudy (cutiecomittee123):

so sin(1)??

OpenStudy (cutiecomittee123):

alright, now I am slightly lost.

OpenStudy (nincompoop):

do not be lost this just requires some practice of fiddling with trig identities and being really familiar with them

OpenStudy (cutiecomittee123):

okay well now what can we do using the sin(1)

OpenStudy (cutiecomittee123):

I am actually stuck now, what can I do?

OpenStudy (anonymous):

google brought up: https://answers.yahoo.com/question/index?qid=20130805043703AAfAth7 which has the solution